# Quantum conditional entropies and fully entangled fraction of states with maximally mixed marginals

### v1

@@ Abstract
The fully entangled fraction (FEF) measures the proximity of a quantum state to maximally entangled states. FEF $>\frac{1}{d}$, in $d\otimes d$ systems is a significant benchmark for various quantum information processing protocols including teleportation. Quantum conditional entropy (QCE) on the other hand is a measure of correlation in quantum systems. Conditional entropies for quantum systems can be negative, marking a departure from conventional classical systems. The negativity of quantum conditional entropies plays a decisive role in tasks like state merging and dense coding. In the present work, we investigate the relation of these two important yardsticks. Our probe is mainly done in the ambit of states with maximally mixed marginals, with a few illustrations from other classes of quantum states. We start our study in two qubit systems, where for the Werner states, we obtain lower bounds to its FEF when the conditional Rényi $\alpha-$entropy is negative. We then obtain relations between FEF and QCE for two qubit Weyl states. Moving on to two qudit states we find a necessary and sufficient condition based on FEF, for the isotropic state to have negative conditional entropy. In two qudit systems the relation between FEF and QCE is probed for the rank deficient and generalized Bell diagonal states. FEF is intricately linked with $k$- copy nonlocality and $k$- copy steerability. The relations between FEF and QCE facilitates to find conditions for $k$- copy nonlocality and $k$- copy steerability based on QCE. We obtain such conditions for certain classes of states in two qubits and two qudits. As a corollary to the relations obtained between QCE and FEF we obtain lower bounds to minimal deterministic work cost for two qubit Werner states and two qudit generalized Bell states.
@@ I Introduction
In the present work, we begin our investigation with $2\otimes 2$ system, wherein we probe the relation of quantum conditional entropies and fully entangled fraction for the Weyl state (states with maximally mixed marginals). In particular, for Werner state, we show that the negativity of conditional von Neumann entropy is a sufficient condition for fully entangled fraction strictly greater than half, and further, we obtain upper bounds to the conditional von Neumann entropy whenever it is useful for teleportation. We observe that fully entangled fraction of two-qubit Werner states has a lower bound in terms of eigenvalues whenever it possesses negative conditional Rényi $\alpha-$entropies. For the two-qubit Weyl state, we show that the negativity of conditional Rényi 2-entropies is a sufficient condition for fully entangled fraction strictly greater than half, and the upper bounds of conditional Rényi 2-entropies are provided if it is useful for teleportation. In the $d\otimes d$ system, for isotropic state [50], we obtain the lower bound of fully entangled fraction whenever its conditional von Neumann entropy is negative. Lower bounds of fully entangled fraction of two-qudit isotropic states is also provided, which guarantees the negativity of conditional Rényi 2-entropies and vice versa. Also, we discuss the relation between fully entangled fraction and conditional entropies for two qudit Werner states, rank deficient states and generalized Bell diagonal states. We then investigate the relation between $k-$copy steerability and conditional entropies of the isotropic state. We find that a two-qubit isotropic state is $k-$copy steerable if conditional von Neumann entropy is negative, and the same holds for conditional Rényi 2-entropies. Moving from $k-$copy steerability to $k-$copy nonlocality, the negativity of conditional Rényi 2-entropy is a sufficient condition for $k-$copy nonlocality of the two-qubit Weyl state. Some two-qubit non-Weyl states show the same behavior under some restriction. We then consider the mixture of arbitrary pure state and maximally mixed noise [39] for dimension $d=3,4,5$ and observe that it is $k-$copy nonlocal whenever they possess negative conditional Rényi 2-entropy and conditional von Neumann entropy. Furthermore, as a corollary to the relations obtained between FEF and QCE, we present an application in the context of work extraction gain. [51].
@@ II Notation and Preliminaries
In this section, we fix the notations, and briefly discuss essential concepts that are important for subsequent study. In this work, we deal with finite-dimensional Hilbert spaces. A quantum system denoted by $X(=A,B)$ is described by finite-dimensional Hilbert space $\mathbb{H}_{X}$. Our work is mainly based on bipartite quantum system $AB=A\otimes B$ that is characterized by Hilbert space $\mathbb{H}_{AB}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$. The Hilbert-Schmidt space $\mathfrak{B}(\mathbb{H}_{AB})$ is the space of bounded linear operators acting on $\mathbb{H}_{AB}$. Quantum states are described by density operators $\rho\in\mathfrak{B}(\mathbb{H}_{AB})$ such that it is positive semi-definite ($\rho\geq 0$), Hermitian (which also follows from positivity) with unit trace. The von Neumann entropy, Rényi $\alpha-$ entropy and Tsallis $\alpha-$ entropy of the quantum state are denoted by $S(.),S_{\alpha}(.),S^{T}_{\alpha}(.)$, respectively. $FEF(.)$ denotes the fully entangled fraction of the quantum state.
@@ II.1 Bloch-Fano Decomposition of density matrices
$\rho_{d_{A}\otimes d_{B}}=\frac{1}{d_{A}d_{B}}[\mathbb{I}_{A}\otimes\mathbb{I}_{B}+\sum_{i=1}^{d_{A}^{2}-1}a_{i}g_{i}^{A}\otimes\mathbb{I}_{B}{}$
$+\sum_{j=1}^{d_{B}^{2}-1}b_{j}\mathbb{I}_{A}\otimes g_{j}^{B}+\sum_{i=1}^{d_{A}^{2}-1}\sum_{j=1}^{d_{B}^{2}-1}t_{ij}g_{i}^{A}\otimes g_{j}^{B}],$ (1)
where $\text{dim}\leavevmode\nobreak\ \mathbb{H}_{A}=d_{A}$ and $\text{dim}\leavevmode\nobreak\ \mathbb{H}_{B}=d_{B}$. The hermitian operators $g_{i}^{k}$ for $k=A,B$ are generalizations of the Pauli matrices, i.e., they are orthogonal $Tr[g_{i}^{k}g_{j}^{k}]=2\delta_{ij}$ and traceless, $Tr[g_{i}^{k}]=0$ and for single qubit systems they coincide with the Pauli matrices. The coefficients $a_{i},b_{j}\in\mathbb{R}$ are the components of the generalized Bloch vectors $\vec{a},\vec{b}$ of the subsystems $A,B$ respectively. The real coefficients $t_{ij}$ are the components of the correlation tensor. For two-qubit systems the density matrices can be represented as,
$\rho_{2\otimes 2}=$ $\frac{1}{4}[\mathbb{I}_{2}\times\mathbb{I}_{2}+\vec{\mathfrak{a}}.\vec{\sigma}\otimes\mathbb{I}_{2}+\mathbb{I}_{2}\otimes\vec{\mathfrak{b}}.\vec{\sigma}+{}$ (2)
$\sum_{i=1}^{3}\sum_{j=1}^{3}t_{ij}\sigma_{i}^{A}\otimes\sigma_{j}^{B}].$
@@
$\rho^{weyl}_{d}=\frac{1}{d^{2}}\left[\mathbb{I}_{A}\otimes\mathbb{I}_{B}+\sum_{i=1}^{d^{2}-1}w_{i}g_{i}^{A}\otimes g_{i}^{B}\right].$ (3)
@@ II.2 Quantum entropies
$S_{\alpha}(\rho_{AB})=\frac{1}{1-\alpha}\log_{2}\left[Tr(\rho_{AB}^{\alpha})\right],\alpha>0,\alpha\neq 1.$ (5)
@@
$S^{T}_{\alpha}(\rho_{AB})=\frac{1}{1-\alpha}[Tr(\rho_{AB}^{\alpha})-1],\alpha>0,\alpha\neq 1.$ (6)
@@ II.3 Fully entangled fraction
In a bipartite quantum system, for any quantum state $\rho_{AB}\in\mathfrak{B}(\mathbb{H}_{AB})$ where $\mathbb{H}_{AB}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$, $\text{dim}\leavevmode\nobreak\ \mathbb{H}_{A}=\text{dim}\leavevmode\nobreak\ \mathbb{H}_{B}=d$, the fully entangled fraction $FEF(\rho)$ is defined as the maximal overlap of quantum state $\rho$ with maximally entangled pure state. In other words, we can also say that it is a measure of proximity of quantum state with maximally entangled pure state. The fully entangled fraction of state $\rho$ is given as
$FEF(\rho_{AB})=\text{max}_{|\phi\rangle\in S}\langle\phi|\rho_{AB}|\phi\rangle.$ (7)
@@
$FEF(\rho_{AB})=\text{max}_{U}\langle\psi_{d}^{+}|U\otimes I\rho U^{\dagger}\otimes I|\psi_{d}^{+}\rangle,$ (8)
@@
$FEF(\rho_{AB})=\frac{1}{4}\left[1+N(\rho_{AB})\right],$ (9)
where $N(\rho)=\text{Tr}|T|$; $|T|=\sqrt{T^{\dagger}T}$ and $T=[t_{ij}]$ is the correlation tensor of the general two qubit density matrix $\rho_{2\otimes 2}$ given by the Equation (2).
@@ III Conditional Entropies and Fully Entangled fraction for two-qubit states
In this section we consider the relationship of the entropic functions like conditional von Neumann entropies and $\alpha$-conditional entropies with fully entangled fraction for different two qubit states like Werner state and Weyl state. Though our results are dependent on the nature of the state, however it gives us a perspective on how the quantifiers of entanglement are connected with entropic measures. As noted earlier, negativity of conditional von Neumann entropy is a benchmark for superdense coding and state merging with quantum advantage. Similarly $FEF>1/d$ for $d\otimes d$ systems decides the efficacy of quantum states in the standard teleportation protocol. Thus the relations between FEF and QCE that we obtain in this and the subsequent sections identify a common class of quantum states which are useful for both teleportation and superdense coding thus opening up the opportunity of using the same resource for multiple tasks.
@@ III.1.1 Relation with conditional von Neumann entropy
$S(A|B)_{\rho^{wer}_{2}(p)}=$ $-3\left(\frac{1-p}{4}\right)\log_{2}\left(\frac{1-p}{4}\right){}$ (15)
$-\left(\frac{1+3p}{4}\right)\log_{2}\left(\frac{1+3p}{4}\right)-1.$
CVNE is negative for $p>0.747614$. The fully entangled fraction of Werner state is given by $FEF(\rho^{wer}_{2})=\frac{1}{4}[1+N(\rho^{wer}_{2})]$, where $N(\rho^{wer}_{2})=Tr|T|$ and $T$ is correlation matrix. So, $FEF(\rho^{wer}_{2})>\frac{1}{2}\Leftrightarrow p>\frac{1}{3}$. Thus, if the Werner state has negative conditional von Neumann entropy (CVNE) then it is useful for teleportation.
@@
If fully entangled fraction of Werner state is strictly greater than half i.e., $FEF(\rho^{wer}_{2})>\frac{1}{2}$, then conditional von Neumann entropy (CVNE) is upper bounded by $3\delta\log_{2}(1/2\delta)$ i.e.,
$FEF(\rho^{wer}_{2})>\frac{1}{2}\Rightarrow\text{CVNE}(\rho^{wer}_{2})<3\delta\log_{2}(1/2\delta),$
where $\delta=\text{min}\{\delta_{1},\delta_{2}\}$, $\delta_{i}\text{ are eigenvalues of }\rho^{wer}_{2}$ and the multiplicity of eigenvalue $\delta_{2}$ is three.
Proof: Let the fully entangled fraction $\text{{\bf{Proof:}} Let the fully entangled fraction }FEF(\rho^{wer}_{2})>\frac{1}{2}{}$
@@
$\Rightarrow-3\left(\frac{1-p}{4}\right)\log_{2}\left(\frac{1-p}{4}\right)-{}$
$\left(\frac{1+3p}{4}\right)\log_{2}\left(\frac{1+3p}{4}\right)-1{}$
@@
$\Rightarrow\text{CVNE}(\rho^{wer}_{2})<3\delta\log_{2}(1/2\delta){}$
@@ III.1.2 Relation with other $\alpha-$ conditional entropies
$S_{\alpha}(A|B)_{\rho^{wer}_{2}(p)}=$ $\frac{1}{1-\alpha}\log_{2}[\frac{(\frac{1+3p}{4})^{\alpha}+3\left(\frac{1-p}{4}\right)^{\alpha}}{2(\frac{1}{2})^{\alpha}}];{}$ (16)
$\alpha>1$
@@
where $\delta=\text{min}\{\delta_{1},\delta_{2}\}$, $\delta_{i}\text{ are eigenvalues of }\rho^{wer}_{2}$ and the multiplicity of eigenvalue $\delta_{2}$ is three.
@@
$\Rightarrow FFE(\rho^{wer}_{2})>\left[\frac{1}{2^{\alpha-1}}-3\delta^{\alpha}\right]^{1/\alpha};\leavevmode\nobreak\ \alpha>1{}$
@@
If conditional Rényi $\alpha-$entropy (CRAE) of two-qubit Werner state is negative, then it is useful for teleportation whenever $\delta<\left[\frac{2^{\alpha}-2^{\alpha-1}}{3.2^{2\alpha-1}}\right]^{\frac{1}{\alpha}}$ where $\delta=\text{min}\{\delta_{1},\delta_{2}\}$, $\delta_{i}\text{ are eigenvalues of }\rho^{wer}_{2}$ and the multiplicity of eigenvalue $\delta_{2}$ is three.
Proof: For $\delta<\left[\frac{2^{\alpha}-2^{\alpha-1}}{3.2^{2\alpha-1}}\right]^{\frac{1}{\alpha}}\Rightarrow\left[\frac{1}{2^{\alpha-1}}-3\delta^{\alpha}\right]^{1/\alpha}>\frac{1}{2}$ and by using Theorem (2) we have $\Rightarrow FFE(\rho^{wer}_{2})>\frac{1}{2}$.

### v2

# On fully entangled fraction and quantum conditional entropies for states with maximally mixed marginals

@@ Abstract
The fully entangled fraction (FEF) measures the proximity of a quantum state to maximally entangled states. FEF $>\frac{1}{d}$, in $d\otimes d$ systems is a significant benchmark for various quantum information processing protocols including teleportation. Quantum conditional entropy (QCE) on the other hand is a measure of correlation in quantum systems. Conditional entropies for quantum systems can be negative, marking a departure from conventional classical systems. The negativity of quantum conditional entropies plays a decisive role in tasks like state merging and dense coding. In the present work, we investigate the relation of these two important yardsticks. Our probe is mainly done in the ambit of states with maximally mixed marginals, with a few illustrations from other classes of quantum states. We start our study in two qubit systems, where for the Werner states, we obtain lower bounds to its FEF when the conditional Rényi $\alpha-$entropy is negative. We then obtain relations between FEF and QCE for two qubit Weyl states. Moving on to two qudit states we find a necessary and sufficient condition based on FEF, for the isotropic state to have negative conditional entropy. In two qudit systems the relation between FEF and QCE is probed for the rank deficient and generalized Bell diagonal states. FEF is intricately linked with $k$- copy nonlocality and $k$- copy steerability. The relations between FEF and QCE facilitates to find conditions for $k$- copy nonlocality and $k$- copy steerability based on QCE. We obtain such conditions for certain classes of states in two qubits and two qudits. Applications of the relations obtained are provided in the context of work extraction, faithful entanglement and entropic uncertainty relations.
Keywords: Quantum conditional entropy, Fully Entangled Fraction, Faithful Entanglement, Steering, Nonlocality
@@ I Introduction
In the present work, we begin our investigation with $2\otimes 2$ system, wherein we probe the relation of quantum conditional entropies and fully entangled fraction for the Weyl state (states with maximally mixed marginals). In particular, for Werner state, we show that the negativity of conditional von Neumann entropy is a sufficient condition for fully entangled fraction strictly greater than half, and further, we obtain upper bounds to the conditional von Neumann entropy whenever it is useful for teleportation. We observe that fully entangled fraction of two-qubit Werner states has a lower bound in terms of eigenvalues whenever it possesses negative conditional Rényi $\alpha-$entropies. For the two-qubit Weyl state, we show that the negativity of conditional Rényi 2-entropies is a sufficient condition for fully entangled fraction strictly greater than half, and the upper bounds of conditional Rényi 2-entropies are provided if it is useful for teleportation. In the $d\otimes d$ system, for isotropic state [50], we obtain the lower bound of fully entangled fraction whenever its conditional von Neumann entropy is negative. Lower bounds of fully entangled fraction of two-qudit isotropic states is also provided, which guarantees the negativity of conditional Rényi 2-entropies and vice versa. Also, we discuss the relation between fully entangled fraction and conditional entropies for two qudit Werner states, rank deficient states and generalized Bell diagonal states. We then investigate the relation between $k-$copy steerability and conditional entropies of the isotropic state. We find that a two-qubit isotropic state is $k-$copy steerable if conditional von Neumann entropy is negative, and the same holds for conditional Rényi 2-entropies. Moving from $k-$copy steerability to $k-$copy nonlocality, the negativity of conditional Rényi 2-entropy is a sufficient condition for $k-$copy nonlocality of the two-qubit Weyl state. Some two-qubit non-Weyl states show the same behavior under some restriction. We then consider the mixture of arbitrary pure state and maximally mixed noise [39] for dimension $d=3,4,5$ and observe that it is $k-$copy nonlocal whenever they possess negative conditional Rényi 2-entropy and conditional von Neumann entropy. Furthermore, as a corollary to the relations obtained between FEF and QCE, we present an application in the context of work extraction [51], faithful entanglement [52, 53, 54], and entropic uncertainty relation [55, 56, 57, 58].
@@ II Notation and Preliminaries
In this section, we fix the notations, and briefly discuss essential concepts that are important for subsequent study. In this work, we deal with finite-dimensional Hilbert spaces. A quantum system denoted by $X(=A,B)$ is described by finite-dimensional Hilbert space $\mathbb{H}_{X}$. Our work is mainly based on bipartite quantum system $AB=A\otimes B$ that is characterized by Hilbert space $\mathbb{H}_{AB}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$. The Hilbert-Schmidt space $\mathfrak{B}(\mathbb{H}_{AB})$ is the space of bounded linear operators acting on $\mathbb{H}_{AB}$. Quantum states are described by density operators $\rho\in\mathfrak{B}(\mathbb{H}_{AB})$ such that it is positive semi-definite ($\rho\geq 0$), Hermitian (which also follows from positivity) with unit trace. The von Neumann entropy, Rényi $\alpha-$ entropy and Tsallis $\alpha-$ entropy of the quantum state are denoted by $S(\cdot),S_{\alpha}(\cdot),S^{T}_{\alpha}(\cdot)$, respectively. $\text{FEF}(\cdot)$ denotes the fully entangled fraction of the quantum state.
@@ II.1 Bloch-Fano Decomposition of density matrices
$\rho_{d_{A}\otimes d_{B}}=\frac{1}{d_{A}d_{B}}[\mathbb{I}_{A}\otimes\mathbb{I}_{B}+\sum_{i=1}^{d_{A}^{2}-1}a_{i}g_{i}^{A}\otimes\mathbb{I}_{B}+\sum_{j=1}^{d_{B}^{2}-1}b_{j}\mathbb{I}_{A}\otimes g_{j}^{B}+\sum_{i=1}^{d_{A}^{2}-1}\sum_{j=1}^{d_{B}^{2}-1}t_{ij}g_{i}^{A}\otimes g_{j}^{B}],$ (1)
where $\text{dim}\leavevmode\nobreak\ \mathbb{H}_{A}=d_{A}$ and $\text{dim}\leavevmode\nobreak\ \mathbb{H}_{B}=d_{B}$. The hermitian operators $g_{i}^{k}$ for $k=A,B$ are generalizations of the Pauli matrices, i.e., they are orthogonal $Tr[g_{i}^{k}g_{j}^{k}]=2\delta_{ij}$ and traceless, $Tr[g_{i}^{k}]=0$ and for single qubit systems they coincide with the Pauli matrices. The coefficients $a_{i},b_{j}\in\mathbb{R}$ are the components of the generalized Bloch vectors $\vec{a},\vec{b}$ of the subsystems $A,B$, respectively. The real coefficients $t_{ij}$ are the components of the correlation tensor. For two-qubit systems the density matrices can be represented as,
$\rho_{2\otimes 2}=\frac{1}{4}[\mathbb{I}_{2}\times\mathbb{I}_{2}+\vec{\mathfrak{a}}.\vec{\sigma}\otimes\mathbb{I}_{2}+\mathbb{I}_{2}\otimes\vec{\mathfrak{b}}.\vec{\sigma}+\sum_{i=1}^{3}\sum_{j=1}^{3}t_{ij}\sigma_{i}^{A}\otimes\sigma_{j}^{B}].$ (2)
@@
$\rho^{weyl}_{d}=\frac{1}{d^{2}}[\mathbb{I}_{A}\otimes\mathbb{I}_{B}+\sum_{i=1}^{d^{2}-1}w_{i}g_{i}^{A}\otimes g_{i}^{B}].$ (3)
@@ II.2 Quantum entropies
$S_{\alpha}(\rho_{AB})=\frac{1}{1-\alpha}\log_{2}\left[Tr(\rho_{AB}^{\alpha})\right],\leavevmode\nobreak\ \alpha>0,\leavevmode\nobreak\ \leavevmode\nobreak\ \alpha\neq 1.$ (5)
@@
$S^{T}_{\alpha}(\rho_{AB})=\frac{1}{1-\alpha}[Tr(\rho_{AB}^{\alpha})-1],\leavevmode\nobreak\ \alpha>0,\leavevmode\nobreak\ \alpha\neq 1.$ (6)
@@ II.3 Fully entangled fraction
In a bipartite quantum system, for any quantum state $\rho_{AB}\in\mathfrak{B}(\mathbb{H}_{AB})$, where $\mathbb{H}_{AB}=\mathbb{H}_{A}\otimes\mathbb{H}_{B}$, $\text{dim}\leavevmode\nobreak\ \mathbb{H}_{A}=\text{dim}\leavevmode\nobreak\ \mathbb{H}_{B}=d$, the fully entangled fraction $\text{FEF}(\rho_{AB})$ is defined as the maximal overlap of quantum state $\rho_{AB}$ with maximally entangled pure state. In other words, we can also say that it is a measure of proximity of quantum state with maximally entangled pure state. The fully entangled fraction of state $\rho$ is given as
$\text{FEF}(\rho_{AB})=\text{max}_{|\phi\rangle\in S}\langle\phi|\rho_{AB}|\phi\rangle,$ (7)
@@
$\text{FEF}(\rho_{AB})=\text{max}_{U}\langle\psi_{d}^{+}|U\otimes I\rho U^{\dagger}\otimes I|\psi_{d}^{+}\rangle,$ (8)
@@
$\text{FEF}(\rho_{AB})=\frac{1}{4}\left[1+N(\rho_{AB})\right],$ (9)
where $N(\rho)=\text{Tr}|T|$; $|T|=\sqrt{T^{\dagger}T}$ and $T=[t_{ij}]$ is the correlation tensor of the general two qubit density matrix $\rho_{AB}$.
@@ III Conditional Entropies and Fully Entangled fraction for two-qubit states
In this section we consider the relationship of the entropic functions like conditional von Neumann entropies and $\alpha$-conditional entropies with fully entangled fraction for different two qubit states like Werner state and Weyl state. Though our results are dependent on the nature of the state, however it gives us a perspective on how the quantifiers of entanglement are connected with entropic measures. As noted earlier, negativity of conditional von Neumann entropy is a benchmark for superdense coding and state merging with quantum advantage. Similarly $\text{FEF}>1/d$ for $d\otimes d$ systems decides the efficacy of quantum states in the standard teleportation protocol. Thus the relations between FEF and QCE that we obtain in this and the subsequent sections identify a common class of quantum states which are useful for both teleportation and superdense coding thus opening up the opportunity of using the same resource for multiple tasks.
@@ III.1.1 Relation with conditional von Neumann entropy
$S(A|B)_{\rho^{wer}_{2}(p)}=-3\left(\frac{1-p}{4}\right)\log_{2}\left(\frac{1-p}{4}\right)-\left(\frac{1+3p}{4}\right)\log_{2}\left(\frac{1+3p}{4}\right)-1.$ (15)
CVNE is negative for $p>0.747614$. The fully entangled fraction of Werner state is given by $\text{FEF}(\rho^{wer}_{2})=\frac{1}{4}[1+N(\rho^{wer}_{2})]$, where $N(\rho^{wer}_{2})=Tr|T|$ and $T$ is correlation matrix. So, $\text{FEF}(\rho^{wer}_{2})>\frac{1}{2}\Leftrightarrow p>\frac{1}{3}$. Thus, if the Werner state has negative conditional von Neumann entropy (CVNE) then it is useful for teleportation.
@@
If fully entangled fraction of Werner state is strictly greater than half i.e., $\text{FEF}(\rho^{wer}_{2})>\frac{1}{2}$, then conditional von Neumann entropy (CVNE) is upper bounded by $3\delta\log_{2}(1/2\delta)$ i.e.,
$\text{FEF}(\rho^{wer}_{2})>\frac{1}{2}\Rightarrow\text{CVNE}(\rho^{wer}_{2})<3\delta\log_{2}(1/2\delta),$
where $\delta=\text{min}\{\delta_{1},\delta_{2}\}$, $\delta_{1}=\frac{1-p}{4}$   and  $\delta_{2}=\frac{1+3p}{4}$ are eigenvalues of $\rho^{wer}_{2}$ with multiplicity $3$ and $1$, respectively.
[…]
@@
$\Rightarrow-3\left(\frac{1-p}{4}\right)\log_{2}\left(\frac{1-p}{4}\right)-\left(\frac{1+3p}{4}\right)\log_{2}\left(\frac{1+3p}{4}\right)-1{}$
@@
$\Rightarrow\text{CVNE}(\rho^{wer}_{2})<3\delta\log_{2}(1/2\delta).{}$
@@ III.1.2 Relation with other $\alpha-$ conditional entropies
$S_{\alpha}(A|B)_{\rho^{wer}_{2}(p)}=$ $\frac{1}{1-\alpha}\log_{2}[\frac{(\frac{1+3p}{4})^{\alpha}+3\left(\frac{1-p}{4}\right)^{\alpha}}{2(\frac{1}{2})^{\alpha}}]\leavevmode\nobreak\ ;\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \alpha>1$ (16)
@@
where $\delta=\text{min}\{\delta_{1},\delta_{2}\}$, $\delta_{1}=\frac{1-p}{4}$   and  $\delta_{2}=\frac{1+3p}{4}$ are eigenvalues of $\rho^{wer}_{2}$ with multiplicity $3$ and $1$, respectively.
@@
$\Rightarrow\text{FEF}(\rho^{wer}_{2})>\left[\frac{1}{2^{\alpha-1}}-3\delta^{\alpha}\right]^{1/\alpha};\leavevmode\nobreak\ \alpha>1.{}$
@@
If conditional Rényi $\alpha-$entropy (CRAE) of two-qubit Werner state is negative, then it is useful for teleportation whenever $\delta<\left[\frac{2^{\alpha}-2^{\alpha-1}}{3.2^{2\alpha-1}}\right]^{\frac{1}{\alpha}}$, where $\delta=\text{min}\{\delta_{1},\delta_{2}\}$, $\delta_{1}=\frac{1-p}{4}$   and  $\delta_{2}=\frac{1+3p}{4}$ are eigenvalues of $\rho^{wer}_{2}$ with multiplicity $3$ and $1$, respectively.
Proof: For $\delta<\left[\frac{2^{\alpha}-2^{\alpha-1}}{3.2^{2\alpha-1}}\right]^{\frac{1}{\alpha}}\Rightarrow\left[\frac{1}{2^{\alpha-1}}-3\delta^{\alpha}\right]^{1/\alpha}>\frac{1}{2}$ and by using Theorem (2) we have $\Rightarrow\text{FEF}(\rho^{wer}_{2})>\frac{1}{2}$.